# Kekulé Superconductivity in Twisted Magic Angle Bilayer Graphene

### v1

@@ Abstract
Motivated by recent scanning tunneling microscopy experiments which report Kekulé ordering in the twisted graphene family, we develop a microscopic theory of superconductivity, specifically for the bi-layer case. This involves an intra-valley, finite-momentum pair-density wave (PDW) which naturally incorporates a Kekulé distortion into the superconducting phase. The PDW state displays three central features: (i) it spontaneously breaks $C_{3}$ rotational symmetry, producing nematic order, and (ii) it possesses a large gap-to-critical-temperature ratio. Moreover, (iii) it yields a quasi-particle density of states having a V-like shape (with a finite value at zero energy) which transitions to a fully gapped, U-shaped spectrum with increasing attraction. These characteristics, including consistency with an allowed Berezinskii-Kosterlitz-Thouless transition, align with key experimental signatures in the twisted graphene family where the U-V transition is observed in the tri-layer case. We find the paired state with a modest interaction strength is near to a BEC-like regime, as appears consistent with the observed extremely short coherence lengths. Together, these results establish a microscopic intra-valley Kekulé PDW state as an essential ingredient towards understanding unconventional superconductivity in twisted graphene.
@@ I Introduction
It is hard to overestimate the excitement which has been generated by the discovery [1] of “high” temperature superconductivity associated with the flat band regime [2] in twisted bi- and tri-layer graphene [3]. Measured so far are transport properties [4, 5], superfluid density[6, 7], tunneling[8, 9, 10], upper critical fields [1, 11], thermodynamical [12] as well as other superconducting characteristics[13, 14, 15, 16].
Among these, scanning tunneling spectroscopy and microscopy [17, 18] have led to remarkable progress in understanding the nature of the correlated insulators. What is most notable is evidence for an inter-valley coherent (IVC) ordering, called Kekulé at $\nu=\pm 2$ and $\pm 3$ which coincides with a $\sqrt{3}\times\sqrt{3}$ reconstruction of the atomic scale unit cell. While there has been much theoretical attention paid to the implications for the correlated insulator regimes[17, 18], a form of Kekulé ordering has also been observed within the superconducting and pseudogap phases in the same interval ($\nu=\pm 2$ to counterpart $\pm 3$.). This, we argue here, is essential to address in order to understand the superconductivity in these twisted graphene, flatband systems.
What has been emphasized in the experimental literature is that (i) the superconductor arises out of a Kekulé order that is qualitatively different from that of the neighboring insulator[17], (ii) that the superconducting and pseudogap phases look similar to one another with, thus far, no readily distinguishable features in imaging [17]. (iii) Importantly, the strength of the IVC order, quantified by the normalized Kekulé peak intensity is maximum at filling factors corresponding to those with very strong pairing, as evidenced by the short coherence lengths and the presence of a pseudogap phase[18].
Figure 1: Cartoon of the twisted bilayer graphene geometry and Cooper Pairs. This serves to contrast intravalley and intervalley pairing. Two graphene Brillouin zones (BZs) are rotated by $\pm\theta$ (gray/red). Two mini-moiré BZs associated with $K/K^{\prime}$ valleys are represented by orange and blue; the $\Gamma$-point is shown in the right mini-BZ. Intravalley pairing involves pairs within each valley and associated mini-BZ, while intervalley pairing involves two different mini-BZs.
Motivated by this STM data, we now consider intra-valley pairing superconductivity in twisted bilayer graphene (TBG), which is a methodology for implementing superconducting Kekulé order [19, 20]. We will see that one can alternatively view this intravalley pairing as a pair density wave (PDW) in which the pairs have non-zero net momentum. In this way, it should be emphasized that “Kekulé superconductivity” as discussed here refers not to a Kekulé pairing mechanism or “glue” but to a Kekulé pairing machinery.
@@
This analysis underlines the importance of systematically addressing, as we do here, the intra-valley pairing case which leads to Kekule superconductivity. This proposal stands in contrast to the majority of current theories, which focus on inter-valley pairing [22, 23, 24, 25, 26] (and its variants [27, 28, 29]). The simple representation in Figure 1 emphasizes the differences in these two scenarios. One cannot, however, rule out that inter-valley and intra-valley pairing may coexist; they originate from distinct, orthogonal symmetry channels of the underlying microscopic attraction.
We emphasize throughout that this intra-valley pairing scenario leads to important and distinct theoretical consequences from the inter-valley case, some of which are rooted in the structure of the flat-band Bloch wavefunctions. The presence of an effect which we term the “quantum textures” presents a notable contrast. Unlike conventional inter-valley pairing where details about the wavefunction are generally absent from the gap equation, this quantum texture, containing wavefunction information, enters directly into the intra-valley gap equation. It strongly modulates the behavior of the order parameter. We will see the interplay between this texture and the fundamental symmetries of a single moiré valley dictates the nature of the resulting superconducting order. We show how this provides direct microscopic insights into the observed numerical phenomenology.
@@ II PDW Order and Microscopic Model
A pair density wave superconductor [30] involves pairs with finite center-of-mass momentum, $2\mathbf{Q}$. The order is bond-centered, forming on the AB/BA sublattice bonds, and in the low-filling regime, the pairing momentum $\mathbf{Q}$ is naturally expected to lie near the Dirac point $\bf K$}. The full order parameter is a time-reversal-symmetric combination of condensates from the two valleys (e.g., $\Delta_{2\mathbf{Q}},\Delta_{-2\mathbf{Q}}]$).
@@
Although intra-valley pairing takes place in two different valleys, it is sufficient for our purposes to analyze a single valley separately. That the two valleys contribute independently is well-justified for small twist angles. Our microscopic theory is based on the Bistritzer-MacDonald (BM) continuum model for TBG, supplemented with a short-range attractive interaction. The BM model captures the essential physics of the moiré superlattice through two interlayer tunneling parameters, $w_{AA}$ and $w_{AB}$. A generic translationally invariant pairing interaction can be written as:
@@
We require pairing interactions to respect all symmetries of the single valley BM model, and involve two different sub-lattices, as we consider AB-bonding order. Among the possible interactions that satisfy these criteria, a natural choice [20, 19] is a nearest-neighbor (NN) attraction within each layer. This is physically motivated by the fact that the interlayer distance is generally larger than the intralayer nearest-neighbor distance.
For definiteness, the numerical results and figures presented throughout this work are based on this NN interaction model. As we will show, the essential properties of the resulting superconducting order originate from the symmetry properties of the BM model. Importantly they are viewed as sufficiently generic and, thus, independent of the details of the assumed short-range attractive interactions.
@@
Here $f_{l}(\mathbf{q})=3^{-1}\sum_{j=0}^{2}\zeta^{jl}\,\exp{-i\mathbf{q}\cdot\boldsymbol{\delta}_{j}}$ , and $\boldsymbol{\delta}_{j}$ represents the nearest-neighbor vectors in honeycomb lattice.
As shown in Figure 1, for intra-valley pairing, Cooper pairs are formed by electrons within the same moiré valley As the mini-Brillouin zone is so tiny, this implies that the interaction is dominated by small momentum transfers, $\mathbf{q}$. Consequently, the phase factor in each $f$ is small [^3] Rigorously, the momentum transfer involved is $\mathbf{q}+\mathbf{G}_{j}$, where $\mathbf{q}$ lies within the first mini-Brillouin Zone and $\mathbf{G}_{j}$ is a moiré reciprocal lattice vector. In the Bistritzer-MacDonald model, $|\mathbf{G}_{j}|$ can be up to three or four times the magnitude of the primitive moiré reciprocal lattice vectors. Nevertheless, the total momentum $|\mathbf{q}+\mathbf{G}_{j}|$ remains significantly smaller than any momentum associated with the atomic scale. $|\mathbf{q}\cdot\boldsymbol{\delta}_{j}|\ll 1$. In this limit, the form factor for the $A$-channel ($l=0$) approaches $f_{0}(\mathbf{q})\to 1$, while the form factors for the chiral channels ($l=1,2$) vanish: $f_{1,2}(\mathbf{q})\to 0$. Thus the dominant superconducting channel corresponds to $A$.  [^4] This conclusion is general and applies to a large class of short-range interaction, as the argument holds if the set of nearest-neighbor vectors, $\boldsymbol{\delta}_{j}$, is replaced by any short-range atomic vector and its $C_{3}$ rotational partners.
Figure 3: Singlet/triplet order parameter as a function of interaction strength $V\equiv V_{0}(q=0)$ near zero temperature $T=0.01$ meV. The filling is taken to be $3/8$ of the lower flat band. Singlet pairs need critical coupling $V_{c}\sim 10$ meV to support a non-zero order parameter, while triplet pairs need $V_{c}\sim 50$ meV. This is due to the different nature of the form factors. Thus we can assume that the superconductivity has singlet order.
@@ III Order Parameter and Quantum Textures
Here the summation is over the band index $m/n$ and momentum $\bf q$} in mini-BZ. We emphasize that in the form factor of Eq. 6, only direct products of Bloch wavefunctions appear. This is to be contrasted with their inner products which would appear in the particle-hole sector. Time-reversal symmetry relates opposite valleys and is absent within a single moiré valley. Consequently, the diagonal form factor $\Lambda_{nn}(\mathbf{q}=0)$ is different from unity, and in practice is often much smaller (as is addressed more explicity in Fig. 2). This is a key distinction between intra-valley pairing and BCS-like inter-valley pairing (or the correlated Kekulé insulator), where the analogue diagonal form factors are close to unity and effectively play no role.
@@
In our numerical simulations, we adopt the isolated flat-band approximation, which depends on the energy separation parameter $E_{\text{band}}$ between the flat and remote bands. In this regime, this parameter and the bandwidth $W$ (which is very small in the absence of strain) satisfy $\Delta/E_{\text{band}}\ll W/\Delta$. This is discussed in more detail in the Methods section. Within this flat band scenario, the pairing physics is governed by a form factor matrix, $\Lambda_{mn}$, where the indices $m,n$ span the flat-band basis. We refer to this matrix as a “quantum texture”, and show it plays a central role here. It encodes the momentum-space structure of the pairing interaction and determines the fundamental properties of the condensate.
Our analysis, based on this flat-band projection, has four important implications for the Kekulé superconductor: (1) The system strongly favors singlet over triplet pairing; this is derived from the $C_{2}\mathcal{T}$ symmetry of the model. (2) The ratio of the zero-temperature gap to the critical temperature is large ($\Delta/T_{c}\gg 2$). (3) The Bogoliubov quasiparticle spectrum spontaneously breaks $C_{3}$ rotational symmetry, indicating an intrinsically nematic superconducting state[34]. (4) The quasi-particle density of states, as measured in tunneling, has a V-like shape (with a “zero bias conductance”) while it transitions to a fully gapped, U-shaped spectrum with increasing attraction.
III.1 (1) Singlet Pairing and $C_{2}\mathcal{T}$-symmetry
We show that the system is more likely to favor singlet than triplet intravalley pairing. As observed in Fig. 3, the coupling required to support singlet superconductivity is substantially smaller than that for the triplet channel. We can understand this numerical finding by demonstrating that it originates from the difference between the form factors in the singlet and triplet channels. This, in turn, derives from the $C_{2}\mathcal{T}$ symmetry of the BM model.
We identify the form factors in each spin (singlet-$s$ and triplet-$t$) channel:
$\Lambda^{s/t}(\mathbf{k},\mathbf{q})=\Lambda(\mathbf{k},\mathbf{q})\pm\Lambda^{\mathsf{T}}(\mathbf{k},-\mathbf{q}).$ (8)
Here the transpose acts on band indices. From the definition of the order parameter in Eq. 7, one sees that the form factor $\Lambda$ together with the coupling $g_{0}$ plays the role of an effective interaction. Thus, the mini-Brillouin-zone (mBZ) average of $|\Lambda^{s/t}(\mathbf{0},\mathbf{q})|$ provides an indicator of which spin channel is favored. Numerically, we confirm that $\langle|\Lambda^{s}(\mathbf{0},\mathbf{q})|\rangle_{\text{mBZ}}>\langle|\Lambda^{t}(\mathbf{0},\mathbf{q})|\rangle_{\text{mBZ}}$. Here $|\cdot|$ denotes the Frobenius norm of the form-factor matrix in the flat-band subspace, and $\langle\cdot\rangle_{\text{mBZ}}$ denotes averaging over the mBZ.
This inequality involving the singlet, triplet states follows from the $C_{2}\mathcal{T}$ symmetry of TBG. To see what this symmetry implies, we consider a toy model consisting of a nondegenerate two-band system which has $C_{2}\mathcal{T}$ symmetry. The model is described by two Bloch functions $\tilde{u}_{\mathbf{G};n}(\mathbf{k})$ with $n=1,2$. Under $C_{2}\mathcal{T}$, one has $\tilde{u}_{\mathbf{G};n}(\mathbf{k})\to\tilde{u}^{*}_{\mathbf{G};n}(\mathbf{k})$,and in this way the Bloch functions can be chosen to be real. Such a two-band toy model is used in Ref. [35] to illustrate how symmetry endows the TBG flat bands with an Euler class. The form factor for Cooper pairing in the toy model is given by:
$\Lambda_{\text{Toy},mn}(\mathbf{Q},\mathbf{q})=\sum_{\mathbf{G}}\tilde{u}_{\mathbf{G};m}(\mathbf{Q}_{+})\,\tilde{u}_{-\mathbf{G};n}(\mathbf{Q}_{-}).$ (9)
This leads to the identity
$\Lambda_{\text{Toy},mn}(\mathbf{Q},\mathbf{q})-\Lambda_{\text{Toy},nm}(\mathbf{Q},-\mathbf{q})=0,$ (10)
and thus the triplet form factor in this toy model vanishes identically. This is a direct consequence of the reality of the Bloch functions.
In the more realistic situation, as for TBG, $C_{2}\mathcal{T}$ acts as $\sigma_{x}\mathcal{K}$ (where $\mathcal{K}$ implies complex conjugation), and consequently Bloch functions are not strictly real. Thus the triplet form factor does not vanish identically. The symmetry $C_{2}\mathcal{T}$ leads to
$\Lambda_{mn,L}(\mathbf{Q},\mathbf{q})\simeq\sum_{\mathbf{G}}u^{*}_{\mathbf{G},LA;m}(\mathbf{Q}_{+})\,u_{-\mathbf{G},LA;n}(\mathbf{Q}_{-}).$ (11)
Here we approximate $f_{0}(\mathbf{q})$ by $f_{0}(0)=1$ since $|f_{0}(\mathbf{q})-f_{0}(0)|\ll 1$. Then with Eq. 11 applied to the triplet state, one sees that the diagonal components of the triplet channel vanish: $\Lambda^{t}_{nn}(\mathbf{Q},\mathbf{q}=\mathbf{0})=0$. This suppression is not found in the singlet channel. This suggests an inequality, quite generally, after averaging over the mBZ $\langle|\Lambda^{s}(\mathbf{Q},\mathbf{q})|\rangle_{\text{mBZ}}>\langle|\Lambda^{t}(\mathbf{Q},\mathbf{q})|\rangle_{\text{mBZ}}$. This inequality then explains the numerical result of Fig. 3 which shows that it takes a much stronger attractive interaction to produce a triplet state.
Figure 4: Temperature dependence of the (singlet) superconducting gap magnitude $|\Delta(T)|$ for interaction strengths $V=-15,-20,-30\,\mathrm{meV}$. The ratios $\Delta/T_{c}$ are approximately $14$, 1010, and $8$, i.e., all large, of order $10$. The normal-state chemical potential is $\mu_{\mathrm{N}}=-0.314\,\mathrm{meV}$. The $T=0$ chemical potentials for the three cases are $\mu=-0.33,-0.38,-0.52\,\mathrm{meV}$, all well above the band bottom at $0.68\,\mathrm{meV}$; thus, the system is not yet in the BEC regime.
III.2 (2) Large $\Delta/T_{c}$ ratio
Solutions of the multiband gap equation show $\Delta/T_{c}\sim 10$, a ratio much larger than the standard BCS value as can be seen from the numerical calculations of Fig. 4. Importantly, we can understand more analytically how this large ratio derives from the properties of the quantum texture. It follows from Figs. 2(c) and (d) that $|\Lambda_{mn}(\mathbf{q})|<1$ for all $\mathbf{q}\in\text{mBZ}$, and indeed $|\Lambda_{mn}(\mathbf{q})|\ll 1$ for many values of $\mathbf{q}$ so that
$\big\langle|\Lambda_{mn}(\mathbf{q})|^{2}\big\rangle_{\text{mBZ}}\ll\big\langle|\Lambda_{mn}(\mathbf{q})|\big\rangle_{\text{mBZ}}\ll 1.$ (12)
Here, $m$ and $n$ refer to the flat-band indices. This inequality reflects the nature of the superconducting order and is a consequence of the facts that (1) the order is located on the AB-bonds, and (2) the condensate in each valley involves only a single moiré valley, which lacks time-reversal and inversion symmetry. (A more detailed discussion of this point is presented in the Methods section.)
To make use of Eq. 12, we consider the gap equation in the simpler case of a single flat band, which is appropriate when $\Lambda_{11}\gg\Lambda_{12}$ and the filling is low. We define $\varphi(\mathbf{q})$ to be the single flat band quantum texture (i.e., $|\Lambda_{11}|$).
In the flat-band limit ($W<2T_{c},~\Delta$) where the bandwidth $W$ is the smallest relevant energy scale in the problem, the gap and $T_{c}$ equations take on a rather unfamiliar form. There is a direct proportionality on the pairing interaction, rather than the exponential dependence found in BCS theory. It can be shown that:
$\Delta\approx\frac{V}{2}\,\big\langle\varphi(\mathbf{q})\big\rangle_{\text{mBZ}},\quad T_{c}\approx\frac{V}{4}\,\big\langle\varphi(\mathbf{q})^{2}\big\rangle_{\text{mBZ}}.$
The ratio $\Delta/T_{c}$ is therefore given by:
$\frac{\Delta}{T_{c}}\approx 2\,\frac{\big\langle\varphi(\mathbf{q})\big\rangle_{\text{mBZ}}}{\big\langle\varphi(\mathbf{q})^{2}\big\rangle_{\text{mBZ}}}.$ (13)
and from Eq. (12) this leads to $\Delta/T_{c}\gg 1$. In particular, for a system with the parameters shown in Fig. 2, we find $\langle\varphi(\mathbf{q})\rangle_{\text{mBZ}}\simeq 0.24$ and $\langle\varphi(\mathbf{q})^{2}\rangle_{\text{mBZ}}\simeq 0.063$. This yields the relations $V/\Delta\sim 10$ and $\Delta/T_{c}\sim 10$, which is consistent the findings in Fig. 4.
In the Methods section we investigate the intermediate bandwidth regime showing that the value of this ratio remains large, approximately $10$, for the same parameter set as is a consequence of these texture effects which lead to the inequality in Eq. (12).

### v2

@@ Abstract
While it has been one of the most important new physics discoveries in the last decade, the nature of superconductivity in the twisted graphene family remains an unsolved problem. Motivated by recent scanning tunneling experiments that report Kekulé ordering in moiré graphene superconductors, we develop a microscopic theory of this superconductivity for the twisted bilayer system. The pairing we find is an intra-valley, finite-momentum pair-density wave (PDW) that intrinsically carries a Kekulé modulation. This state exhibits four salient features: (i) spontaneous breaking of $C_{3}$ rotation symmetry, producing nematic order (ii)with triplet pairing; and (iii) a quasiparticle density of states that evolves from a V-shaped profile to a fully gapped, U-shaped spectrum as the attraction increases which is accompanied by (iv) systematic behavior of the temperature dependent zero bias conductance. These features align with key experimental signatures. We find, as well, that with only modest interaction strengths, the state is near to a BEC-like phase, consistent with the observed extremely short coherence lengths. Taken together, these results identify a microscopic intra-valley Kekulé PDW as a compelling candidate for unconventional superconductivity in the twisted graphene family.
@@ I Introduction
It is hard to overestimate the excitement which has been generated by the discovery [Cao2018a] of “high” temperature superconductivity associated with the flat band regime [Bistritzer2011] in twisted bi- and tri-layer graphene [Nuckolls2024]. Measured so far are transport properties [Polshyn2019, Cao2020], superfluid density[Banerjee2025, Tanaka2025], tunneling[Jiang2019, Oh2021, Xie2019], upper critical fields [Cao2018a, Park2021], thermodynamical [Zondiner2020] as well as other superconducting characteristics[Yankowitz2019, Sharpe2019, Stepanov2019, Saito2019].
Among these, scanning tunneling spectroscopy and microscopy [Nuckolls2023, Kim2023] have led to progress in understanding the nature of the correlated insulators. What is most remarkable is that evidence for an inter-valley coherent (IVC) ordering, called Kekulé was anticipated theoretically [Wagner2022, Kwan2021]. This appears at $\nu=\pm 2$ and $\pm 3$ which coincides with a $\sqrt{3}\times\sqrt{3}$ reconstruction of the atomic scale unit cell. While there has been much theoretical attention [Wagner2022, Kwan2021] paid to the implications for the correlated insulator regimes[Nuckolls2023, Kim2023], a form of Kekulé ordering has also been observed within the superconducting and pseudogap phases in the same interval ($\nu=\pm 2$ to counterpart $\pm 3$.). This, we argue here, is essential to address in order to understand the superconductivity in these twisted graphene, flatband systems.
What has been emphasized in the experimental literature is that (i) the superconductor arises out of a Kekulé order that is qualitatively different from that of the neighboring insulator[Nuckolls2023], (ii) that the superconducting and pseudogap phases look similar to one another with, thus far, no readily distinguishable features in imaging [Nuckolls2023]. (iii) Importantly, the strength of the IVC order, quantified by the normalized Kekulé peak intensity is maximum at filling factors corresponding to those with very strong pairing, as evidenced by the short coherence lengths and the presence of a pseudogap phase[Kim2023].
Figure 1: Cartoon of the twisted bilayer graphene geometry and Cooper Pairs. This serves to contrast intravalley and intervalley pairing. Two graphene Brillouin zones (BZs) are rotated by $\pm\theta$ (gray/red). Two mini-moiré BZs associated with $K/K^{\prime}$ valleys are represented by orange and blue; the $\Gamma/M$-point is shown in the right/left mini-BZ. Intravalley pairing involves pairs within each valley and associated mini-BZ, while intervalley pairing involves two different mini-BZs.
Motivated by this STM data, we now consider intra-valley pairing superconductivity in twisted bilayer graphene (TBG), which is a methodology for implementing superconducting Kekulé order [Roy2010, Tsuchiya2016, yafis2025]. We will see that one can alternatively view this intravalley pairing as a pair density wave (PDW) in which the pairs have non-zero net momentum. In this way, it should be emphasized that “Kekulé superconductivity” as discussed here refers not to a Kekulé pairing mechanism or “glue” but to a Kekulé pairing machinery.
@@
This analysis underlines the importance of systematically addressing, as we do here, the intra-valley pairing case which leads to Kekule superconductivity. This proposal stands in contrast to the majority of current theories, which focus on inter-valley pairing [You2019, Wu2018, Lian2019, Cea2021, Hu2019, Julku2020, Christos2023] (and its variants [Khalaf2021, Gonzalez2019, Po2018]). The simple representation in Figure 1 emphasizes the differences in these two scenarios. One cannot, however, rule out that inter-valley and intra-valley pairing may coexist; they originate from distinct, orthogonal symmetry channels of the underlying microscopic attraction.
We emphasize throughout that this intra-valley pairing scenario leads to important and distinct theoretical consequences from the inter-valley case, some of which are rooted in the structure of the flat-band Bloch wavefunctions. The presence of an effect which we term the “quantum textures” presents a notable contrast. Unlike conventional inter-valley pairing where details about the wavefunction are generally absent from the gap equation, this quantum texture, containing wavefunction information and multi-band effects, enters directly into the intra-valley gap equation. It strongly modulates the behavior of the order parameter. We will see the interplay between this texture and the fundamental symmetries of a single moiré valley dictates the nature of the resulting superconducting order.
@@ II PDW Order and Microscopic Model
A pair density wave superconductor [Agterberg2020] involves pairs with finite center-of-mass momentum, $2\mathbf{Q}$. The order is bond-centered, forming on the AB/BA sublattice bonds, and in the low-filling regime, the pairing momentum $\mathbf{Q}$ is naturally expected to lie near the Dirac point $\bf K$}. The full order parameter is a time-reversal-symmetric combination of condensates from the two valleys (e.g., $\Delta_{2\mathbf{Q}},\Delta_{-2\mathbf{Q}}]$).
@@
Although intra-valley pairing takes place in two different valleys, it is sufficient for our purposes to analyze a single valley separately. That the two valleys contribute independently is well-justified for small twist angles. Our microscopic theory is based on the Bistritzer-MacDonald (BM) continuum model for TBG[Bistritzer2011, PhysRevB.100.035448, PhysRevB.99.035111], supplemented with a short-range attractive interaction. The BM model captures the essential physics of the moiré superlattice through two interlayer tunneling parameters, $w_{AA}$ and $w_{AB}$. A generic translationally invariant pairing interaction can be written as:
@@
We require pairing interactions to respect all symmetries of the single valley BM model, and involve two different sub-lattices, as we consider AB-bonding order. Among the possible interactions that satisfy these criteria, a natural choice [Roy2010, Tsuchiya2016] is a nearest-neighbor (NN) attraction within each layer. This is physically motivated by the fact that the interlayer distance is generally larger than the intralayer nearest-neighbor distance.
For definiteness, the numerical results and figures presented throughout this work are based on this NN interaction model. As we will show, the essential properties of the resulting superconducting order originate from the symmetry properties of the BM model. Importantly they are viewed as generally generic and, thus, independent of the details of the assumed short-range attractive interactions.
@@
Here $f_{l}(\mathbf{q})=3^{-1}\sum_{j=0}^{2}\zeta^{jl}\,\exp{-i\mathbf{q}\cdot\boldsymbol{\delta}_{j}}$ , and $\boldsymbol{\delta}_{j}$ represents the nearest-neighbor vectors in the honeycomb lattice.
As shown in Figure 1, for intra-valley pairing, Cooper pairs are formed by electrons within the same moiré valley As the mini-Brillouin zone is little, this implies that the interaction is dominated by small momentum transfers, $\mathbf{q}$. Consequently, the phase factor in each $f$ is small [^3] Rigorously, the momentum transfer involved is $\mathbf{q}+\mathbf{G}_{j}$, where $\mathbf{q}$ lies within the first mini-Brillouin Zone and $\mathbf{G}_{j}$ is a moiré reciprocal lattice vector. In the Bistritzer-MacDonald model, $|\mathbf{G}_{j}|$ can be up to three or four times the magnitude of the primitive moiré reciprocal lattice vectors. Nevertheless, the total momentum $|\mathbf{q}+\mathbf{G}_{j}|$ remains significantly smaller than any momentum associated with the atomic scale. $|\mathbf{q}\cdot\boldsymbol{\delta}_{j}|\ll 1$. In this limit, the form factor for the $A$-channel ($l=0$) approaches $f_{0}(\mathbf{q})\to 1$, while the form factors for the chiral channels ($l=1,2$) vanish: $f_{1,2}(\mathbf{q})\to 0$. Thus the dominant superconducting channel corresponds to $A$  [^4] This conclusion is general and applies to a large class of short-range interaction, as the argument holds if the set of nearest-neighbor vectors, $\boldsymbol{\delta}_{j}$, is replaced by any short-range atomic vector and its $C_{3}$ rotational partners. .
Figure 3: The figure shows the grand-canonical thermodynamic potential $\Omega$ of the singlet and triplet pairing states as a function of interaction strength $V\equiv V_{0}(q=0)$. Results are computed at near-zero temperature $T=0.001~\mathrm{meV}$ and chemical potential $\mu=-0.59~\mathrm{meV}$. The plotted quantity is $\Omega-\Omega_{N}$, measured relative to the normal-state value $\Omega_{N}$. The inset shows the corresponding superconducting order-parameter magnitude $|\Delta|$ for the two states. For the interaction strengths shown, the triplet state has a lower $\Omega$ than the singlet state, indicating that triplet pairing is thermodynamically favored.
@@ III Order Parameter and Quantum Textures
Here the summation is over the band index $m/n$ and momentum $\bf q$} in the mini-BZ. We emphasize that in the form factor of Eq. 6, only direct products of Bloch wavefunctions appear. This is to be contrasted with their inner products which would appear in the particle-hole sector. Time-reversal symmetry relates opposite valleys and is absent within a single moiré valley. Consequently, the diagonal form factor $\Lambda_{nn}(\mathbf{Q},\mathbf{q}=0)$ generally deviates from unity, and the off-diagonal components $\Lambda_{mn}$ acquire sizeable magnitudes that are central to the pairing machinery. This structure, which reflects the unique quantum texture of the intra-valley sector, is addressed more explicitly in Fig. 2. This is a key distinction between intra-valley pairing and BCS-like inter-valley pairing (or the correlated Kekulé insulator), where the analogue diagonal form factors are close to unity.
In our numerical simulations, we perform a truncation of those remote bands in the vicinity of the flat-bands. This analysis confirms that a two-isolated- flat- band approximation is adequate, largely because of the sizeable separation $E_{\text{band}}$ between the flat and remote bands. Here we have that the bandwidth $W$ (which is very small in the absence of strain) satisfies $\Delta/E_{\text{band}}\ll W/\Delta$. More precision can be obtained when more bands are added, as is discussed in the Methods section.
@@
The pairing physics is governed by a form factor matrix, $\Lambda_{mn}$, where the indices $m,n$ span the band basis. We refer to this matrix as a “quantum texture” which plays a central role in our analysis as it reflects the momentum and band space structure of the pairing interaction, thus determining the fundamental properties of the condensate.
III.1 Establishing the Thermodynamically Stable Phase
For a given pairing momentum $\bf Q$}, one can solve the gap equation self-consistently to find possible solutions. Because not all solutions represent thermodynamically stable states it is necessary to investigate their stability by establishing that they correspond to a global minimum of the grand-canonical potential $\Omega$ at a given chemical potential $\mu$.
In the PDW state under consideration, the thermodynamic potential is a function $\Omega(\mu,{\bf Q},\Delta_{L},{\bf S})$, where we label the layer index $L=1,2$ and $\mathbf{S}$ represents the pair spin. There are two branches of solutions satisfying the gap equation: $\Delta_{1}=\Delta_{2}$ and $\Delta_{1}=-\Delta_{2}$. Both branches respect mirror symmetry. The first branch, which is the physical solution, corresponds to a global minimum for a fixed $\bf Q$} and $\mu$, while the second branch represents a saddle point. Thus, throughout the paper we use $\Delta\equiv\Delta_{1}$ to denote the order-parameter magnitude.
In this stable ground state with $\Delta_{1}=\Delta_{2}$, $\bf Q$} is a variational parameter which is constrained to be within the mini-BZ  [^5] The zero-current condition is automatically satisfied once the time-reversal symmetric partner at the other valley is included. This differs from the known FF state, where $Q$ must be solved self-consistently to satisfy the zero-current condition. Important here also is that a conventional FF state is generally associated with a saddle point (see Wang et al, Phys. Rev. B 97, 134513), which is circumvented in multilayer systems where there are more degrees of freedom to enable stability.. For a given $\mu$ and spin configuration (singlet or triplet), the potential $\Omega$ is minimal at ${\bf Q}=M$ in the mini-BZ. This indicates that the slow moiré-scale modulation observed in STM data must be determined by the momentum $M$ relative to Dirac points.
We note that the relative magnitude of $\Omega$ for the singlet and triplet states depends on the pairing momentum $\bf Q$}. Importantly, when $\bf Q$} is chosen to be the $M$-point, the triplet state has a lower $\Omega$ than the singlet, as is consistent with what is observed experimentally [Cao2021]. This is shown in Fig. 3, where it should be clear that the energy scales associated with the strength of the needed attraction are of the order of meV. This should be quite accessible and it is notably smaller than estimates [Julku2020] of the order of $0.1$ eV, based on inter-valley pairing. This can be seen in more detail in Fig. 4.
The difference between the singlet and triplet pairings [^6] Since the Hamiltonian preserves full spin rotation symmetry, we select the triplet channel with $S_{z}=0$ without loss of generality. is encoded in the structure of the form factors:
$\Lambda^{s/t}(\mathbf{Q},\mathbf{q})=\frac{1}{2}\left[\Lambda(\mathbf{Q},\mathbf{q})\pm\Lambda^{\mathsf{T}}(\mathbf{Q},-\mathbf{q})\right].$ (8)
Here, $s$ and $t$ denote singlet and triplet, respectively, and the transpose acts on the band indices. Due to $C_{2}\mathcal{T}$ symmetry, the diagonal component of $\Lambda^{t}$ (triplet channel) is strongly suppressed compared to the singlet channel [^7] The $C_{2}\mathcal{T}$ symmetry enforces the condition $\Lambda^{t}_{nn}(\mathbf{Q},\mathbf{q}=\mathbf{0})=0$.. However, there are two active flat bands near the chemical potential. As in Ref. [Christos2023] which considered a different pairing configuration, the off-diagonal pairing plays a crucial additional role here regarding the competition between singlet and triplet states which depend on $\bf Q$}. To summarize, the stable order is characterized by two equal layer condensates, pairing momentum at $M$, and spin-triplet pairing.